OpenStudy (anonymous):

Find the product of (x + 3)(x − 3).

OpenStudy (diyadiya):

\[(a+b)(a-b)=a^2-b^2\]

OpenStudy (diyadiya):

\[(x)^2-(3)^2=x^2-9\]

OpenStudy (anonymous):

thank you! (: